a man is 56 years old and his is 24 years old in how many years the father will be twice as old as his son at that time?
Answer by stanbon(75887)

a man is 56 years old and his son is 24 years old in how many years the father will be twice as old as his son at that time?
-------
Equation:
Let the number of years be "x":
-----
56 + x = 2(24 + x)
----
56 + x = 48 + 2x
x = 8 years (# of years till father is twice son's age)
-------------------
Cheers,
Stan H.